# 1–1–1 Tennodai, Tsukuba-shi, 305–8573, Ibaraki, Japan

### v1

@@ 1.1 Background
Recent advances in information and communication technologies have enabled consumers to browse and purchase a wide variety of items through online services. Meanwhile, consumers often struggle to find items that match their preferences among a plethora of options. To overcome this information overload problem, recommender systems [1, 2] have been widely implemented in online services. These systems generate personalized lists of unknown but potentially preferred items. In fact, integrating these systems into online platforms simplifies the search process and improves user experience [24].
@@ 3 Robust portfolio optimization model
Specifically, for each $u\in U$ we consider the following ranges of variation in the expectation and covariance of ratings:
@@
where $\Gamma^{(\mu)},\Gamma^{(\sigma)}\in\mathbb{Z}_{+}$ are integer-valued parameters for limiting the cardinality of $S^{(\mu)}$ and $S^{(\sigma)}$.
@@ 4 Experiments
In this section, we evaluate the effectiveness of our robust portfolio optimization method for recommendation through numerical experiments.
@@ 4.1 Datasets
The Yahoo! R3 dataset is pre-divided into training and testing sets, which contain 15,400 and 5,400 users, respectively. Each user rated 1,000 different songs on a scale of 1 to 5, and the testing set contained 10 ratings per user. We extracted ratings from the 5,400 users included in both the training and testing sets.
@@ 4.2 Experimental setup
For each user $u\in U$, the candidate item set $I_{u}$ was composed of the items rated by the user in the corresponding testing set. The number of recommended items was set to $N=10$ for the MovieLens 100K dataset and $N=5$ for the Yahoo! R3 dataset. We set $\alpha=0.2$ as the hyperparameter for adjusting the mean–variance tradeoff. We performed shrinkage estimation of rating covariance matrices as
@@
$\delta_{ui}^{(\mu)}=\frac{\sqrt{\hat{\sigma}_{ii}}}{\sqrt{\max\{1,n_{i}\}}}\quad((u,i)\in U\times I),$
$\delta_{ij}^{(\sigma)}=\frac{0.2\cdot\hat{\sigma}_{ij}}{\sqrt{\max\{1,n_{ij}\}}}\quad((i,j)\in I\times I),$
@@ 4.4 Results for the MovieLens 100K dataset
These results on the MovieLens 100K dataset (Fig. 1) demonstrate that our robust portfolio optimization method can significantly improve the accuracy or diversity of recommendations by appropriately tuning the cardinality parameters $\Gamma^{(\mu)}$ and $\Gamma^{(\sigma)}$, compared to not considering the uncertainty in the expectation and covariance of ratings (i.e., $\Gamma^{(\mu)}=\Gamma^{(\sigma)}=0$).
@@ 4.5 Results for the Yahoo! R3 dataset
These results on the Yahoo! R3 dataset (Fig. 2) demonstrate that our robust portfolio optimization method can improve recommendation diversity without reducing recommendation accuracy, compared to not considering the uncertainty in the expectation and covariance of ratings (i.e., $\Gamma^{(\mu)}=\Gamma^{(\sigma)}=0$). This positive result can be attributed to the fact that the accuracy of rating prediction was lower on the Yahoo! R3 dataset (RMSE: 1.42) than on the MovieLens 100K dataset (RMSE: 0.92), as mentioned in Section 4.2.

### v2

@@ 1.1 Background
Recent advances in information and communication technologies have enabled consumers to browse and purchase a wide variety of items through online services. Meanwhile, consumers often struggle to find items that match their preferences among a plethora of options. To overcome this information overload problem, recommender systems [1] have been widely implemented in online services. These systems generate personalized lists of unknown but potentially preferred items. In fact, integrating these systems into online platforms simplifies the search process and improves user experience [24].
@@ 3 Robust portfolio optimization model
Specifically, for each user $u\in U$ we consider the following ranges of variation in the expectation and covariance of ratings:
@@
where $\Gamma^{(\mu)},\Gamma^{(\sigma)}\in\mathbb{Z}_{+}$ are integer-valued parameters for limiting the cardinalities of $S^{(\mu)}$ and $S^{(\sigma)}$.
@@ 4 Experiments
In this section, we evaluate the effectiveness of our robust portfolio optimization method for recommendation through numerical experiments. All experiments were performed on a Mac OS 14.4 computer equipped with an Apple M3 processor @ 4.05 GHz (8 cores) and 24 GB RAM.
@@ 4.1 Datasets
The Yahoo! R3 dataset is pre-divided into training and testing sets, which contain 15,400 and 5,400 users, respectively. Each user rated some of the 1,000 songs on a scale of 1 to 5, and the testing set contained 10 ratings per user. We extracted ratings from the 5,400 users included in both the training and testing sets.
@@ 4.2 Experimental setup
For each user $u\in U$, the candidate item set $I_{u}$ was composed of the items rated by the user in the corresponding testing set. The number of recommended items was set to $N=10$ for the MovieLens 100K dataset and $N=5$ for the Yahoo! R3 dataset. We set $\alpha=0.2$ as the hyperparameter for adjusting the mean–variance trade-off. We performed shrinkage estimation of rating covariance matrices as
@@
$\delta_{ui}^{(\mu)}=\frac{\sqrt{\hat{\sigma}_{ii}}}{\sqrt{\max\{1,n_{i}\}}}\quad((u,i)\in U\times I),\quad\delta_{ij}^{(\sigma)}=\frac{0.2\cdot\hat{\sigma}_{ij}}{\sqrt{\max\{1,n_{ij}\}}}\quad((i,j)\in I\times I),$
@@ 4.4 Results for the MovieLens 100K dataset
These results on the MovieLens 100K dataset (Fig. 1) demonstrate that our robust portfolio optimization method can significantly improve the accuracy or diversity of recommendations by appropriately tuning the cardinality parameters $\Gamma^{(\mu)}$ and $\Gamma^{(\sigma)}$, compared to not considering the uncertainty in the expectation or covariance of ratings (i.e., $\Gamma^{(\mu)}=\Gamma^{(\sigma)}=0$).
@@ 4.5 Results for the Yahoo! R3 dataset
These results on the Yahoo! R3 dataset (Fig. 2) demonstrate that our robust portfolio optimization method can improve recommendation diversity without reducing recommendation accuracy, compared to not considering the uncertainty in the expectation or covariance of ratings (i.e., $\Gamma^{(\mu)}=\Gamma^{(\sigma)}=0$). This positive result can be attributed to the fact that the accuracy of rating prediction was lower on the Yahoo! R3 dataset (RMSE: 1.42) than on the MovieLens 100K dataset (RMSE: 0.92), as mentioned in Section 4.2.
4.6 Discussion
Yasumoto and Takano [37] demonstrated that increasing $\alpha$ in Eq. (2) improves recommendation diversity. This trend is similar to our results when increasing $\Gamma^{(\sigma)}$. This similarity likely arises because both $\alpha$ and $\Gamma^{(\sigma)}$ encourage the inclusion of item pairs with low covariance in the recommendation list.
The previous studies [32, 33] in information retrieval demonstrated that mean–variance portfolio analysis increases diversity of ranked documents and improves retrieval performance. Thus, our robust optimization model is also potentially effective in ranking documents. Moreover, our recommendation performance can be further improved by properly tuning $\alpha$.
Finally, we focus on experimental results of robust optimization models in other domains, such as personalized pricing [16] and coupon allocation [31]. These studies demonstrated that the model performance and its standard error tended to improve as the degree of considered uncertainty raised. A similar trend was observed in our experiments, where the F1 scores and their standard errors improved with increasing $\Gamma^{(\mu)}$ on both real-world datasets.